home / study / business / economics / economics questions and answers / in an imaginary economy, consumers buy only sandwiches and magazines. the fixed basket consists ... Question: In an imaginary economy, consumers buy only sandwiches and magazines. The fixed basket consists o... In an imaginary economy, consumers buy only sandwiches and magazines. The fixed basket consists of 25 sandwiches and 40 magazines. in 2006, a sandwich cost $4.50 and a magazine cost $3.99. In 2007, a sandwich cost $6.75. If the inflation rate in 2007 was 21 percent, then how much did a magazine cost in 2007?
Answer #1

basket cost in 2006

=25*4.50+40*3.99

=272.100

basket cost in 2007

=272.10*(1+21%)

=329.241

basket cost in 2007=329.241

25*6.75+40*magazine cost=329.241

magazine cost=(329.241-25*6.75)/40=4.01

the above is answer..